OpenStudy (anonymous):

Simplify this radical expression. Use absolute value symbols when needed.... Answer Choices: A) -k^3 b) k^3 C) k^5 D) -k^5

OpenStudy (anonymous):

\[\sqrt[5]{-k^15}\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt[5]{-k^{15}}\] \[\Large \sqrt[5]{-1*k^{15}}\] \[\Large \sqrt[5]{-1}*\sqrt[5]{k^{3*5}}\] \[\Large \sqrt[5]{-1}*\sqrt[5]{\left(k^3\right)^5}\] I'll let you finish

jhonyy9 (jhonyy9):

@owlphant95 do you understand it ?

OpenStudy (anonymous):

Not really...

jimthompson5910 (jim_thompson5910):

Are you familiar with the idea that \[\Large \sqrt[5]{x^5} = x\] ?

jhonyy9 (jhonyy9):

@owlphant95 do you know the property of exponents ?

jhonyy9 (jhonyy9):

like ((x)^a)^b =(x)^(a*b)

jhonyy9 (jhonyy9):

for example

jhonyy9 (jhonyy9):

@owlphant95 so what not is clearly here if you have learned the property of exponents this is very easy

OpenStudy (anonymous):

Hello i am sorry i dont understand it at all

jhonyy9 (jhonyy9):

hello are you here ?